OpenStudy (anonymous):

Tickets=19, People =200.Option 1= $2.00 Admission $.50 Per ticket ,19x$.50x 200 people = $1900 ticket Profit +400$= 2300 together. Option 2= No Admission, 1.00$ per ticket ,$1.00x19x200=3800. Profit Equal income(the money take in) minus expenses. for each option, Write an Equation to find the profit Y of selling X tickets. Assume that 200 people will attend and use $1000 as your expenses. Graph the equtions in the same Coordinate Plane. For which numbers of tickets sold are the two options equal? what is the profit or loss for this number of tickets? also what would the Formula be for

OpenStudy (anonymous):

hey Saifoo could you help?

OpenStudy (saifoo.khan):

Hello Spence, im trying. ^_^

OpenStudy (anonymous):

its something to do with Linear Equations

OpenStudy (saifoo.khan):

we have to make 2 equations i think....

OpenStudy (anonymous):

Correct one for Option 1 and one for Option 2 then ive got to take that Equation and Graph it. IM stuck on making the Equation

OpenStudy (saifoo.khan):

is that the whole original question?

OpenStudy (anonymous):

Umm if you want i could Give you the Web Site that the Project is on?

OpenStudy (saifoo.khan):

Okay. mind emailing the info? or will u giv it here?

OpenStudy (anonymous):

ill send it here its easyer ittle just be a link

OpenStudy (anonymous):

Saifoo, what the heck?

OpenStudy (saifoo.khan):

OK. @mario, lol. mathteacher helped u already.

OpenStudy (anonymous):

that wasnt my problem

OpenStudy (anonymous):

That should be the link and Mario please dont start any thing on my MAth PRoblem i need to get this done

OpenStudy (anonymous):

http://openstudy.com/#/updates/4edeb1cde4b05ed8401ab5ac this is ^^^

OpenStudy (saifoo.khan):

mario, i will be there after this.

OpenStudy (anonymous):

and sorry. i need to get mine done too.

OpenStudy (anonymous):

its okay :), this is a Project and Due Tomarrow :/ SaiFoo if your on the Link im on Question 2.) . if you need the Information from 1 ill be happy to give it

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

just tell me if you need the other Information

OpenStudy (saifoo.khan):

p = profit, you'd get p = 200(2) + 0.50x - 1000 and p = 1.00x - 1000

OpenStudy (anonymous):

is that the Equation? how would i Graph that though?

OpenStudy (saifoo.khan):

im trying to fig t out.

OpenStudy (anonymous):

okay thanks

OpenStudy (saifoo.khan):

OPTION A: p = 200(2) + 0.50x - 1000 p = 0.5x -600 Option B: p = 1.00x - 1000

OpenStudy (saifoo.khan):

im still not sure about these.

OpenStudy (anonymous):

may i ask were we get the -600 at?

OpenStudy (saifoo.khan):

it says 200 people. and the entry fee is 2. so, 2(200) now we are going to subtract this from our main expenses. 1000-400

OpenStudy (anonymous):

ahh okay , i see it now thanks

OpenStudy (saifoo.khan):

uw

OpenStudy (anonymous):

so we pretty much have the Equation part down ??

OpenStudy (saifoo.khan):

yes.

OpenStudy (anonymous):

how do we Graph this?

OpenStudy (saifoo.khan):

i will brb

OpenStudy (saifoo.khan):

rewrite that: opt A y = 0.5x -600 Option B: y = 1.00x - 1000

OpenStudy (saifoo.khan):

Now when x-intercept put y = 0 And when y intercept out x = 0

OpenStudy (anonymous):

wait how do i Graph it >.<

OpenStudy (anonymous):

so (0,?) and (?,0)